solve dx/2x(y+z^2) = dy/y(2y+z^2) = dz/z^3 find two solutions


1
Expert's answer
2021-04-13T23:30:17-0400

The given equation is incorrect, The correct equation is-


"\dfrac{dx}{2x(y+z^2)} = \dfrac{dy}{2y(y+z^2)} = \dfrac{dz}{z^3}"


Taking first two terms-

"\dfrac{dx}{2x(y+z^2)}=\dfrac{dy}{2y(y+z^2)}"


"\Rightarrow \dfrac{dx}{x}=\dfrac{dy}{y}"


Integrating both the sides-

"logx=logy+logc_1\\

x=yc_1\\c_1=\dfrac{x}{y}~~~~~-(1)"


Taking Last two terms we have-

"\dfrac{dy}{2y(y+z^2)}=\dfrac{dz}{z^3}"


"\Rightarrow z^3dy-2y(y+z^2)dz=0"


Divide both sides by "-z^3y^2"


"-\dfrac{1}{y^2}\dfrac{dy}{dz}+\dfrac{2}{zy}=\dfrac{-2}{z^3}"


Let "\dfrac{1}{y}=t\implies \dfrac{-1}{y^2}\dfrac{dy}{dz}=\dfrac{dt}{dz}"


The above equation becomes-


"\dfrac{dy}{dz}+\dfrac{2t}{z}=\dfrac{-2}{z^3}"


On comparing it with Linear differential equation, we get-

"P=\dfrac{2}{z},Q=\dfrac{-2}{z^3}"


Integrated factor, I.F.="e^{\int P.dz}

 =e^{\int \dfrac{2}{z}dz}

 =e^{2logz}=z^2"

Its solution is-


"t\times I.F.=\int I.F.\times Q dz"


"\Rightarrow t(z^2)=\int z^2\times \dfrac{-2}{z^3}dz"


"\Rightarrow t\times z^2=-2logz+c_2"


"\Rightarrow c_2=\dfrac{z^2}{y}+2logz"


The solution of the given equation is-

"\phi(c_1,c_2)=0\\

\phi(\dfrac{x}{y},\dfrac{z^2}{y}+2logz)=0"